OpenStudy (anonymous):

For each equation, choose the statement that describes its solution. If applicable, give the solution. 4(w-2)-3w = 2(w-6) no solution w=? All real numbers are solutions 4(y+1)-y = 3(y-1)+3 no solution y=? All real numbers are solutions

OpenStudy (anonymous):

The first one would be 4w-8-3w=2w-12, w-8=2w-12, 4=w,w=4 The second one is 4y+4-y=3y-3+3, 3y+4=3y, 3=0, no solution